OpenStudy (anonymous):

The sum of 3 consecutive integral numbers is 117. Find the numbers.

OpenStudy (anonymous):

x + (x+1) + (x+2) = 117

hartnn (hartnn):

let smallest integer be x so 2 integers are x,x+1,x+2 sum = 117 can u form an equation now ?

OpenStudy (anonymous):

x is the first number in the sequence

OpenStudy (anonymous):

oh heck guess and check!

OpenStudy (anonymous):

jeez not equations lets try 40 40+41+42=123 no too big try again

hartnn (hartnn):

guess,by dividing by 3

OpenStudy (anonymous):

don't forget you are a thinking person, not an equation solving machine

OpenStudy (anonymous):

But thinking gives equation right

hartnn (hartnn):

is rcb getting any of this ? :P

OpenStudy (anonymous):

38+39+40=117

OpenStudy (anonymous):

And YES DIVIDING BY THREE GIVES MIDDLE VALUE.

hartnn (hartnn):

thought so

OpenStudy (anonymous):

Because, (x-1)+x+(x+1)=117 { where x is the middle value} 3x=117